# SDFs from Unoriented Point Clouds using Neural Variational Heat Distances

### v1

@@ 3.1.2 Far Field Normals
To obtain a good approximation of surfaces with highly curved features, we use small time steps $\tau$ in the variational heat flow time step 4 to obtain a consistent field of normals $n^{\tau}$. However, we observed that for such small time steps, the heat solution $u^{\tau}$ has very small gradient norms $\lvert\nabla u^{\tau}\rvert$ further away from the surface $\mathcal{S}$. This leads to a lack of robustness when normalizing $\nabla u^{\tau}$ to compute $n^{\tau}$. To mitigate this, we compute a second solution $u^{\hat{\tau}}$ of the variational heat flow problem 4 for a significantly larger time step size $\hat{\tau}$ and use the resulting less precise normal field $n^{\hat{\tau}}$ in the far field. To still obtain an overall smooth normal field, we blend the gradient fields of the heat solution around a critical value $\kappa$ of the original solution. Hence, we define the blended normal field
@@
for $\beta_{\kappa}(s)=\mu(\tfrac{s}{\kappa})$ using the $C^{1}$ smooth, cubic Hermite blending function $\mu\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$ with $\mu(s)=1$ for $s<0$, $\mu(s)=\tfrac{1}{4}(2s+1)(2s-2)^{2}$ for $0\leq s\leq 1$ and $\mu(s)=0$ for $s>1$. In fact, for $u^{\tau}\geq\kappa$ we have $n^{\ast}(x)=n^{\tau}(x)$ and in the limit for decreasing values of $u^{\tau}$ we obtain $n^{\ast}(x)=n^{\hat{\tau}}(x)$. We simply use $n^{\ast}$ instead of $n^{\tau}$ in the smoothed normal approximation objective (10). When minimizing ${\mathcal{E}_{n}}$ for this normal field, we combine an accurate signed distance in the near field and a proper identification of the zero-level set with a robust extension of the signed distance in the far field. In the applications we choose $\kappa=\tfrac{3}{5}\max_{u^{\tau}}$, where $\max_{u^{\tau}}$ is the maximum of $|u^{\tau}|$ on the cell centers of a regular grid with grid size $0.0375$ (cf. section 3.1.2).
The results in fig. 6 indeed show that using the blended normal field leads to much better gradient norms and level sets farther from the surface. However, they degrade near the medial axis of the shapes. Thus creating a trade-off. If one is only interested in accurate results in a narrow band, the near field is sufficient. For our quantitative evaluations in the comparison section, we have always used the far field approach.
@@ 3.2.1 Inside & Outside
[…]
@@ 3.2.2 Smoothed Characteristic Functions
The normal approximation objective ${\mathcal{E}_{n}}$ depends on the sets $[\phi<0]$ and $[\phi>0]$, which means that it depends on the function $\phi$ in a non-smooth manner due to the thresholding operation. To facilitate the numerical minimization of ${\mathcal{E}_{n}}$, we introduce a smooth approximation of the corresponding characteristic functions such that the objective becomes differentiable (see fig. 8). To this end, we define $\eta_{\delta}(s)\coloneqq\eta(\delta^{-1}s)$ using the $C^{1}$ smooth, cubic Hermite blending function $\eta\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$ with $\eta(s)=1$ for $s<-1$, $\eta(s)=\tfrac{1}{4}(s+2)(s-1)^{2}$ for $|s|\leq 1$ and $\eta(s)=0$ for $s>1$. For $\delta\to 0$ and $s\neq 0$, $\eta_{\delta}(s)$ converges to $\chi_{\{s<0\}}$ and thus $\eta_{\delta}(\phi(\cdot))$ converges pointwise to $\chi_{[\phi<0]}$ on $[\phi\neq 0]$ for a fixed $\phi$. Finally, this allows us to introduce the smoothed approximation
@@ 4 Experimental Results
We implemented our method using the PyTorch framework. To represent $u$ and $\phi$, we use fully connected neural networks with periodic activation functions as proposed by [30]. Concretely, we use four hidden layers of size 256 and the sine function as activation in all experiments. We use the Adam optimizer [18] to approximately solve our minimization problems. All experiments were trained for a total number of 50 epochs, each consisting of 1000 batches of constant batch size. During training, we decrease the learning rate up to $10^{-8}$ using the PyTorch function ReduceLROnPlateau with patience $2$; as an initial learning rate, we choose $10^{-4}$. All experiments took under an hour, including both optimization steps. For all experiments, we choose $\tau=0.005$ for the small time step and $\hat{\tau}=0.1$. We plan to release the code.
@@ 4.1 Comparisons
We compare with 1-Lip [6] that focuses on having low eikonal errors and good distance approximations by making use of a Lipschitz network architecture. It additionally requires oriented normals. Moreover, we consider HESS [34], SALD [1], and HotSpot [33]. For details on the models and point clouds, see appendix B.
As evaluation metrics, we use two metrics defined on the input point cloud, evaluating the SDF and its gradient on the zero-level set, and two metrics defined in a narrow band sampled near the surface, evaluating the SDF and the eikonal error:
  1. $\mathbf{E_{recon}^{\mathcal{S}}}$
    The squared $L^{2}$-error of the learned SDFs on sampled ground truth surface points to measure the reconstruction error.
  2. $\mathbf{E_{recon}^{n}}$
    The $L^{1}$-cosine distance between the gradients of the SDF on the zero-level set and the normals of the mesh.
  3. $\mathbf{E_{SDF}\;\>}$
    The $L^{1}$-error between the learned SDF and the ground truth signed distance to the input surface on a narrow band around the surface.
  4. $\mathbf{E_{eik}}\ \ \;\>$
    The median eikonal error on a narrow band around the surface.
See section 4.2 for more details.
@@ 4. $\mathbf{E_{eik}}\ \ \;\>$
In fig. 9, we show a quantitative comparison of the different methods over a range of shapes (cf. appendix B, fig. 15). We plot the surface reconstruction error $\mathbf{E_{recon}^{\mathcal{S}}}$ against the two measures for SDF quality $\mathbf{E_{eik}}$ and $\mathbf{E_{SDF}}$. The results show that our method achieves a balanced trade-off between an accurate surface reconstruction and a good SDF approximation. This is in contrast to HESS, which produces good reconstruction results but has larger eikonal and SDF errors and 1-Lip, which has low eikonal errors but loses detail in the surface reconstruction. Some results of SALD and HESS had flipped signs, which we inverted before measuring the errors.
@@
Figure 10 shows a qualitative comparison between the zero-level sets extracted with Marching Cubes for the different methods. Moreover, we plot for points, sampled in a narrow band near the surface, the neural distances compared the ground truth signed distance to the mesh. While HESS, HotSpot, and our method all achieve an accurate shape reconstruction, our method leads to the most accurate SDF estimation. We also note that HESS has a very accurate SDF estimation only very close to the surface and $1$-Lip consistently underestimates the distance, as expected from these methods. Those observations are also reflected in the corresponding table.
@@
Our quadrature for the surface integral 12 accounting for the input point cloud density allows us to compute accurate SDFs even for spatially varying point densities. To demonstrate this, we consider a capped torus shape that can be described as the zero-level of a closed-form SDF [27]. We consider a dense, a sparse, and a non-uniform point cloud as well as an example with noise added to the surface; see table 1 for a visualization of the point clouds and quantitative values comparing the uniform and non-uniform case. All methods can deal with noisy and sparse input data. We provide the quantitative evaluation in the appendix table 3. However, HotSpot and SALD have difficulties dealing with the highly non-uniform point cloud, mixing inside and outside of the shape. In contrast, our SDF error only changes by a magnitude of $10^{-4}$. In fig. 11, we show scatter plots for the SDF evaluation on the non-uniform capped torus.
@@
Figure 9: Quantitative evaluation for different methods applied to various models (cf. appendix B, fig. 15), illustrating that our method effectively balances surface reconstruction and SDF approximation; top: eikonal error $\mathbf{E_{eik}}$ against surface error $\mathbf{E_{recon}^{\mathcal{S}}}$; bottom: SDF error $\mathbf{E_{SDF}}$ against surface error $\mathbf{E_{recon}^{\mathcal{S}}}$.
@@
Figure 11: SDF error scatter plots for the non-uniformly sampled capped torus from table 1. Only 1-Lip and our method show a globally high consistency of the SDF. All other methods significantly deviate from the ground truth in the narrow band.
@@
4.2 Parameters and Error Metrics
Weighting of the loss terms.
We have a single weighting parameter $\lambda_{\mathrm{fit}}$ in front of ${\mathcal{E}_{\mathrm{fit}}}$, constraining the SDF to be close to zero on the input point cloud. We see a trade-off between a good surface fitting minimizing ${\mathcal{E}_{\mathrm{fit}}}$ and the normal alignment loss ${\mathcal{E}_{n}}$. This is reflected in the decreasing reconstruction error for increasing $\lambda_{\mathrm{fit}}$ which comes with an increase of the eikonal error, see fig. 12. Extremely high values of $\lambda_{\mathrm{fit}}$ might lead to surfaces with small artifacts. However, for the range covered in the plot, all surfaces can be recovered without artifacts. In all our experiments, we choose $\lambda_{\mathrm{fit}}=100$ as our focus lies on good SDF qualities.
Steepness of blending function.
We use a smoothed blending function $\eta_{\delta}$ to approximate the characteristic functions of the sets $[\phi<0]$ and $[\phi>0]$ in (14). We study the dependence of our result on the parameter $\delta$, which corresponds to the steepness of $\eta_{\delta}$. For $\delta\in[0.001,0.01]$ the $L^{2}$ surface error only changes by orders of $10^{-6}$ and the eikonal error in a narrow band by $10^{-4}$, see table 2. Hence, the choice of this parameter is not very sensitive. In all our experiments, we choose $\delta=0.005$.
Evaluation metrics.
using a set $\mathcal{M}$ of $50$k distinct on-surface points sampled on the ground truth surface. To this end, we exploit that all our considered models come with a ground truth mesh. Further, we evaluate the normal alignment error on the input surface by computing the cosine distance:
where $n(x)$ denote the discrete normals of the respective ground truth mesh at the triangle centers $x$ and $\mathcal{F}$ the set of face centers of the respective mesh. We further evaluate the quality of the results on a set $\mathcal{N}$ of $10$k points in a narrow band with distances in $[-0.1,0.1]$ from $\mathcal{S}$. The set $\mathcal{N}$ was generated once per model using rejection sampling based on the ground truth SDF and was subsequently used consistently across all evaluations. We compute the SDF error compared to a ground truth signed distance $\mathrm{d}(\cdot,\mathcal{S})$ to the mesh surface $\mathcal{S}$:
as well as the eikonal error:
@@ 4.2 Parameters and Error Metrics
4.3 Application: Neural PDE Heat Flow on Neural Surfaces
In this section, we derive a level set method for the geometric heat equation following [3, 13] and formulate it as a minimizing movement scheme to make it accessible for a neural network method (cf. [26]).
@@ 4.3 Application: Neural PDE Heat Flow on Neural Surfaces
${\mathcal{E}^{\phi}_{\mathrm{MM}}}(w)$ $\coloneqq\int_{\Omega}(w-w^{k})^{2}+\tau\left|P\nabla w\right|^{2}\,\mathrm{d}x$
@@
The minimization problem in (18) decouples over all level sets $\mathcal{S}_{c}$. This is only guaranteed for this continuous formulation. A numerical approximation comes with only an approximate decoupling. In practice, we confine to a narrow band using a smooth blending function $\mu_{\sigma}$ with $\mu_{\sigma}(s)=1$ on $[-\tfrac{\sigma}{2},\tfrac{\sigma}{2}]$ and vanishes outside $(-\sigma,\sigma)$, where in our application $\sigma=0.05$.
Altogether, a time-discrete, fully implicit Euler scheme can be formulated as a variational problem and thus solved numerically using a neural network approach. Figure 13 (bottom) shows the results of neural heat diffusion on the neural zero-level set of the Max Planck head, in comparison to a piecewise affine, continuous finite element method on the original triangulation (top) from which the point cloud was extracted. Additionally, we present results of the finite element heat flow computation on the mesh extracted from the neural zero-level set (middle). We observe that the three approaches lead to very similar results qualitatively, indicating that our neural-PDE-on-neural-surface framework yields valid results without requiring to extract the mesh of the neural level set.

### v2

@@ 3.1.2 Far Field Normals
To obtain a good approximation of surfaces with highly curved features, we use small time steps $\tau$ in the variational heat flow time step ˜4 to obtain a consistent field of normals $n^{\tau}$. However, we observed that for such small time steps, the heat solution $u^{\tau}$ has very small gradient norms $\lvert\nabla u^{\tau}\rvert$ further away from the surface $\mathcal{S}$. This leads to a lack of robustness when normalizing $\nabla u^{\tau}$ to compute $n^{\tau}$. To mitigate this, we compute a second solution $u^{\hat{\tau}}$ of the variational heat flow problem ˜4 for a significantly larger time step size $\hat{\tau}$ and use the resulting less precise normal field $n^{\hat{\tau}}$ in the far field. To still obtain an overall smooth normal field, we blend the gradient fields of the heat solution around a critical value $\kappa$ of the solution $u^{\tau}$. Hence, we define the blended normal field
@@
for $\beta_{\kappa}(s)=\mu(\tfrac{s}{\kappa})$ using the $C^{1}$ smooth, cubic Hermite blending function $\mu\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$ with $\mu(s)=1$ for $s<0$, $\mu(s)=\tfrac{1}{4}(2s+1)(2s-2)^{2}$ for $0\leq s\leq 1$ and $\mu(s)=0$ for $s>1$. In fact, for $u^{\tau}\geq\kappa$ we have $n^{\ast}(x)=n^{\tau}(x)$ and in the limit for decreasing values of $u^{\tau}$ we obtain $n^{\ast}(x)=n^{\hat{\tau}}(x)$. We simply use $n^{\ast}$ instead of $n^{\tau}$ in the smoothed normal approximation objective (10). When minimizing $\mathcal{E}_{n}$} for this normal field, we combine an accurate signed distance in the near field and a proper identification of the zero-level set with a robust extension of the signed distance in the far field. In the applications we choose $\kappa=\tfrac{3}{5}\max_{u^{\tau}}$, where $\max_{u^{\tau}}$ is the maximum of $|u^{\tau}|$ on the cell centers of the regular grid used below to distinguish inside and outside of $\mathcal{S}$ (cf. Section˜3.2).
The results in Figure˜7 indeed show that using the blended normal field leads to much better gradient norms and level sets farther from the surface. However, they degrade near the medial axis of the shapes. Thus creating a trade-off. If one is only interested in accurate results in a narrow band, the near field is sufficient. For our quantitative evaluations in the comparison section, we have always used the far field approach.
@@ 3.2.1 Inside & Outside
We construct the sets $\mathcal{B}^{-}$ on the inside of $\mathcal{S}$ and $\mathcal{B}^{+}$ on the outside using the following simple and efficient algorithm (see Figure˜8). We consider a regular grid with edge length $h>0$ and cell centers $\alpha h$ with $\alpha\in\mathbb{Z}^{3}$ and denote the corresponding open cells by $B_{\alpha}$. Step 1. We mark all cells whose closure has a non-empty intersection with $\mathcal{S}$ (i.e., containing points of the given point cloud) as interfacial cells. Step 2. Beginning with one cell intersecting the boundary of $\Omega$ that is not marked as interfacial, we successively mark each of the six neighboring cells for every outside cell $B_{\alpha}$, which are not yet marked, as outside. Step 3. We mark all remaining cells as interior. Finally, we define $\mathcal{B}^{+}$ as the union of all outside cells and $\mathcal{B}^{-}$ as the union of all inside cells. In practice, a rather coarse grid is sufficient, e.g., we used $64^{3}$ grid points and grid size $h=0.0375$. We note that for a very sparse point cloud, we may require a larger $h$. For example, for the sparse bean model in Table˜1, we used $2h$ as the grid size, which was the only case where different parameters were needed.
@@ 3.2.2 Smoothed Characteristic Functions
The normal approximation objective $\mathcal{E}_{n}$} depends on the sets $\phi<0]$ and $\phi>0]$, which means that it depends on the function $\phi$ in a non-smooth manner due to the thresholding operation. To facilitate the numerical minimization of $\mathcal{E}_{n}$}, we introduce a smooth approximation of the corresponding characteristic functions such that the objective becomes differentiable (see Figure˜9). To this end, we define $\eta_{\delta}(s)\coloneqq\eta(\delta^{-1}s)$ using the $C^{1}$ smooth, cubic Hermite blending function $\eta\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$ with $\eta(s)=1$ for s<1s<-1, $\eta(s)=\tfrac{1}{4}(s+2)(s-1)^{2}$ for $|s|\leq 1$ and $\eta(s)=0$ for $s>1$. For $\delta\to 0$ and $s\neq 0$, $\eta_{\delta}(s)$ converges to $\chi_{\{s<0\}}$ and thus $\eta_{\delta}(\phi(\cdot))$ converges pointwise to $\chi_{[\phi<0]}$ on $\phi\neq 0]$ for a fixed $\phi$. Finally, this allows us to introduce the smoothed approximation
@@ 4 Experimental Results
We implemented our method using the PyTorch framework. All experiments took under an hour, including both optimization steps. The code is available at https://github.com/sweidemaier/HeatSDF
4.1 Parameters and Error Metrics
Network parameters.
To represent $u$ and $\phi$, we use fully connected neural networks with periodic activation functions as proposed by [32]. Concretely, we use four hidden layers of size 256 and the sine function as activation in all experiments. We use the Adam optimizer [19] to approximately solve our minimization problems. All experiments were trained for a total number of 50 epochs, each consisting of 1000 batches of constant batch size. During training, we decrease the learning rate up to $10^{-8}$ using the PyTorch function ReduceLROnPlateau with patience $2$; as an initial learning rate, we choose $10^{-4}$.
Weighting of the loss terms.
We have a single weighting parameter $\lambda_{\mathrm{fit}}$ in front of $\mathcal{E}_{\mathrm{fit}}$}, constraining the SDF to be close to zero on the input point cloud. We see a trade-off between a good surface fitting minimizing $\mathcal{E}_{\mathrm{fit}}$} and the normal alignment loss $\mathcal{E}_{n}$}. This is reflected in the decreasing reconstruction error for increasing $\lambda_{\mathrm{fit}}$ which comes with an increase of the eikonal error, see Figure˜10. Extremely high values of $\lambda_{\mathrm{fit}}$ might lead to surfaces with small artifacts. However, for the range covered in the plot, all surfaces can be recovered without artifacts. In all our experiments, we choose $\lambda_{\mathrm{fit}}=100$ as our focus lies on good SDF qualities.
Timestep and blending function.
For all experiments, we choose $\tau=0.005$ for the small time step and $\hat{\tau}=0.1$. We use a smoothed blending function $\eta_{\delta}$ to approximate the characteristic functions of the sets $\phi<0]$ and $\phi>0]$ in (14). We study the dependence of our result on the parameter $\delta$, which corresponds to the steepness of $\eta_{\delta}$. For $\delta\in[0.001,0.01]$ the $L^{2}$ surface error only changes by orders of $10^{-6}$ and the eikonal error in a narrow band by $10^{-4}$, see Table˜2. Hence, the choice of this parameter is not very sensitive. In all our experiments, we choose $\delta=0.005$.
Evaluation metrics.
We use two metrics defined on the input point cloud, evaluating the SDF and its gradient on the zero-level set, and two metrics defined in a narrow band sampled near the surface, evaluating the SDF and the eikonal error.
using a set $\mathcal{M}$ of 5050k distinct on-surface points sampled on the ground truth surface. To this end, we exploit that all our considered models come with a ground truth mesh. Further, we evaluate the normal alignment error on the input surface by computing the cosine distance:
where $n(x)$ denote the discrete normals of the respective ground truth mesh at the triangle centers $x$ and $\mathcal{F}$ the set of face centers of the respective mesh.
We further evaluate the quality of the results on a set $\mathcal{N}$ of 1010k points in a narrow band with distances in [0.1,0.1][-0.1,0.1] from $\mathcal{S}$. The set $\mathcal{N}$ was generated once per model using rejection sampling based on the ground truth SDF and was subsequently used consistently across all evaluations. We compute the SDF error compared to a ground truth signed distance $\mathrm{d}(\cdot,\mathcal{S})$ to the mesh surface $\mathcal{S}$:
as well as the eikonal error:
@@ 4.2 Comparisons
We compare with 1-Lip [6] that focuses on having low eikonal errors and good distance approximations by making use of a Lipschitz network architecture. It additionally requires oriented normals. Moreover, we consider HESS [36], SALD [1], and HotSpot [35]. For details on the models and point clouds, see Appendix˜B.
@@
In Figure˜11, we show a quantitative comparison of the different methods over a range of shapes (cf. Appendix˜B, Figure˜17). We plot the surface reconstruction error $\mathbf{E_{recon}^{\mathcal{S}}}$ against the two measures for SDF quality $\mathbf{E_{eik}}$ and $\mathbf{E_{SDF}}$. The results show that our method achieves a balanced trade-off between an accurate surface reconstruction and a good SDF approximation. This is in contrast to HESS, which produces good reconstruction results but has larger eikonal and SDF errors and 1-Lip, which has low eikonal errors but loses detail in the surface reconstruction. Some results of SALD and HESS had flipped signs, which we inverted before measuring the errors.
@@
Figure˜12 shows a qualitative comparison between the zero-level sets extracted with Marching Cubes for the different methods. Moreover, we plot for points, sampled in a narrow band near the surface, the neural distances compared the ground truth signed distance to the mesh. While HESS, HotSpot, and our method all achieve an accurate shape reconstruction, our method leads to the most accurate SDF estimation. We also note that HESS has a very accurate SDF estimation only very close to the surface and $1$-Lip consistently underestimates the distance, as expected from these methods. Those observations are also reflected in the corresponding table.
@@
Our quadrature for the surface integral ˜12 accounting for the input point cloud density allows us to compute accurate SDFs even for spatially varying point densities. To demonstrate this, we consider a capped torus shape that can be described as the zero-level of a closed-form SDF [29]. We consider a dense, a sparse, and a non-uniform point cloud as well as an example with noise added to the surface; see Table˜1 for a visualization of the point clouds and quantitative values comparing the uniform and non-uniform case. All methods can deal with noisy and sparse input data. We provide the quantitative evaluation in the appendix Table˜3. However, HotSpot and SALD have difficulties dealing with the highly non-uniform point cloud, mixing inside and outside of the shape. In contrast, our SDF error only changes by a magnitude of $10^{-4}$. In Figure˜13, we show scatter plots for the SDF evaluation on the non-uniform capped torus.
@@
Figure 11: Quantitative evaluation for different methods applied to various models (cf. Appendix˜B, Figure˜17), illustrating that our method effectively balances surface reconstruction and SDF approximation; top: eikonal error $\mathbf{E_{eik}}$ against surface error $\mathbf{E_{recon}^{\mathcal{S}}}$; bottom: SDF error $\mathbf{E_{SDF}}$ against surface error $\mathbf{E_{recon}^{\mathcal{S}}}$.
@@
Figure 13: SDF error scatter plots for the non-uniformly sampled capped torus from Table˜1. Only 1-Lip and our method show a globally high consistency of the SDF. All other methods significantly deviate from the ground truth in the narrow band.
@@
4.3 Application: Neural PDE Heat Flow on Neural Surfaces
In this section, we demonstrate that our learned SDF is sufficiently accurate to solve a PDE directly on the neural field using a level set method. Following [3, 13], we start by deriving a level set method for the geometric heat equation and formulate it as a minimizing movement scheme to make it accessible for a neural network-based approach (cf. [28]).
@@ 4.3 Application: Neural PDE Heat Flow on Neural Surfaces
${\mathcal{E}^{\phi}_{\mathrm{MM}}}[w]$ $\coloneqq\int_{\Omega}(w-w^{k})^{2}+\tau\left|P\nabla w\right|^{2}\,\mathrm{d}x$
@@
The minimization problem in (18) decouples over all level sets $\mathcal{S}_{c}$. This is only guaranteed for this continuous formulation. A numerical approximation comes with only an approximate decoupling. In practice, we confine to a narrow band using a smooth blending function $\mu_{\sigma}$ with $\mu_{\sigma}(s)=1$ on $\tfrac{\sigma}{2},\tfrac{\sigma}{2}]$ and vanishes outside $(-\sigma,\sigma)$, where in our application $\sigma=0.05$.
Altogether, a time-discrete, fully implicit Euler scheme can be formulated as a variational problem and thus solved numerically using a neural network approach. Figure˜14 (bottom) shows the results of neural heat diffusion on the neural zero-level set of the Max Planck head, in comparison to a piecewise affine, continuous finite element method on the original triangulation (top) from which the point cloud was extracted. Additionally, we present results of the finite element heat flow computation on the mesh extracted from the neural zero-level set (middle). We observe that the three approaches lead to very similar results qualitatively, indicating that our neural-PDE-on-neural-surface framework yields valid results without requiring to extract the mesh of the neural level set.